OpenStudy (anonymous):

(a)An angle measures 60° . What is the measure of its supplement? (b)An angle measures 46° . What is the measure of its complement?

OpenStudy (cwrw238):

the supplement of an angle A = 180 - A complement is 90 - A

OpenStudy (anonymous):

how you get it

OpenStudy (cwrw238):

replace A by 60 in 180-A to get answer to (a) replace A by 46 in 90 -A to get answer to (b)

OpenStudy (cwrw238):

so (a) is 180 - 60 = ?

OpenStudy (anonymous):

120

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

what about (b)?

OpenStudy (anonymous):

90-46=44?

OpenStudy (cwrw238):

right!

OpenStudy (anonymous):

thannks!!

OpenStudy (cwrw238):

yw